# SvANet: A Scale-variant Attention-based Network for Small Medical Object Segmentation

### v3

@@ IV-C Results for the Dataset for Only Ultra-small Objects
IV-D Ablation Studies
IV-D1 Main Components Ablation
Ablation studies were conducted and discussed in this section to investigate the influence of each module of SvANet. The results of ablation studies for the MCBottleneck, MCAttn, cross-scale guidance, SvAttn, and AssemFormer modules are presented in Tab. IV. As indicated in the first and second columns of Tab. IV, the inclusion of MCBottleneck and MCAttn results in improvements of +0.15% & +0.52% mDice and +0.40% & +0.22% sensitivity, respectively, suggesting that both modules significantly contribute to enhancing the diagnosis and accuracy of positive cases. Furthermore, when equipped with cross-scale guidance and SvAttn, SvANet achieves an additional +0.41% & +0.24% in mDice and +0.53% & +0.51% in sensitivity. The ablation of AssemFormer leads to an increase of +0.16% in mDice and +0.67% in sensitivity. The p-value for mDice is below 0.05, confirming the result reliability. In summary, MCAttn and cross-scale guidance are critical modules for enhancing prediction accuracy, while AssemFormer and MCBottleneck are vital in improving the accuracy of positive diagnostics in medical object segmentation.
IV-D2 MCAttn vs. Other Advanced Attention Methods
IV-E Negative Case Studies
Examples of visualization results for ultra-small and small medical objects using the top three performance methods, HRNet, TransNetR, and SvANet in the ATLAS, KiTS23, and SpermHealth dataset are presented in Fig. 5. As indicated by the first two rows in Fig. 5, it is possible to overlook overlapped medical objects, particularly the ultra-small ones. For example, both HRNet and TransNetR failed to identify an ultra-small tumor inside the liver and missed an ultra-small cyst at the edge of the kidney. Although HRNet and TransNetR identified the kidney with a complete boundary, they incorrectly emphasized holes that are not part of the kidney in the example image. However, SvANet accurately differentiated between organs and their infected regions. Besides, SvANet reflected the morphological details of the liver, hepatic tumor, kidney, and cyst in the example image, aligning closely with the ground truth annotations. The lower-right image in Fig. 5 shows that SvANet precisely located all sperms’ positions and effectively recognized the region of the broken-line tail of an abnormal sperm. Conversely, tested methods like HRNet struggled to distinguish the broken-line shape. Moreover, TransNetR misclassified an unhealthy sperm head as healthy, as indicated by a green subregion in Fig. 5d. These visualization results align with the findings discussed in Sec. IV-B and Sec. IV-C, suggesting that SvANet holds significant potential for application in general small medical object recognition across various medical imaging modalities.
V Conclusion
This paper introduces SvANet, a novel network to improve the segmentation of small medical objects, aiding in the detection of life-threatening diseases and supporting in vitro fertilization. The experimental outcomes demonstrate that the SvANet is highly effective in distinguishing medical objects of various sizes. SvANet consistently outperformed other SOTA methods, achieving 96.12%, 96.11%, 89.79%, 84.15%, 80.25%, 73.05%, and 72.58% in mDice for segmenting objects with less than 1% image area across multiple datasets, including KiTS23, ISIC 2018, ATLAS, PolypGen, TissueNet, FIVES, and SpermHealth. Additionally, the visualization results confirm that SvANet accurately identifies the locations and morphologies of all medical objects, demonstrating its exceptional capability in the segmentation of small medical objects. These findings underscore the potential of SvANet as a significant advancement in medical imaging.
Supplement
Area Computation
The same instance across different images is treated as an independent object because several datasets do not have instance labels during inference. For a fair comparison, each object instance is counted once per image. Area computation results indicate that 18% of retinal vasculatures in FIVES [1], 40% of skin lesions in ISIC 2018 [2], 63% of polyps in PolypGen [3], 74% of hepatic tumors and livers in ATLAS [6], 99% of kidneys and corresponding neopasms in KiTS23 [4], and 100% tissue cells in TissueNet [7]. The area calculation codes are publically available at https://github.com/anthonyweidai/circle_extractor.
Network Instantiation
Inference Time
@@ V Conclusion
This section quantifies the inference characteristics of the tested networks, including the number of parameters (# Parameters), multiply-accumulate operations (MACs), and inferencing speed. The number of classes was set to eight, and other configurations were consistent with those described in Sec. IV-A2. The inference speed results, averaged over 1000 runs, are presented in Tab. VI. Tab. VI illustrates that SvANet achieved a real-time analysis of medical images with 46.91 frames per second (FPS), which is twice as fast as UNet and 19% faster than TransNetR. This performance level is suitable for self-examination and clinical diagnostic applications.
TABLE VI: Model size and inference time comparison.
@@
AssemFormer Module Understanding
To better understand AssemFormer’s capabilities, we visualized the feature maps generated by AssemFormer, as shown in Fig. 6. In Fig. 6, progressing from left to right, the AssemFormer increasingly highlights smaller areas that more accurately align with the ground truth, especially notable in scenarios with fewer medical objects. For instance, the first row of Fig. 2 demonstrates how the thin lines of retinal vasculature and light reflections are initially emphasized, becoming progressively thicker. Subsequently, these lines become shorter and more focused on a smaller region corresponding to the optic disc location, as depicted in the second first of Fig. 2ef. Large-scale distortions, such as noise or compression defects, play a role in this observed trend where the concentration of feature maps intensifies with a deeper layer. The pattern of increased feature map concentration is consistent across the segmentation of various medical objects, including skin lesions, polyps, hepatic tumors, livers, kidneys, tissue cells, and sperms.
Figure 6: Visualization of output feature maps from AssemFormer. Columns (a) and (b) display the original images and their corresponding ground truth annotations, respectively. Columns (c), (d), and (f) illustrate the output feature maps from AssemFormer at stages 1, 2, 3, and 4 across seven benchmark datasets.
Robustness analysis
To quantify the robustness and adaptability of SvANet vs. other SOTA methods, receiver operating characteristic (ROC) curves of tested methods in the SpermHealth dataset are employed and illustrated in Fig. 7. The ROC curve of SvANet, represented by the red line in Fig. 7, blends nearest towards the top-left corner, with the highest area under the curve (AUC) of 0.9905, surpassing other SOTA methods by up to AUC of +0.0995. Besides, the ROC curve of HSNet is close to the lower-right corner and under all other curves, with the lowest AUC of 0.8910. The ROC and AUC results of HSNet are consistent with Tab. III, demonstrating that HSNet struggled to recognize sperms.
Full Negative Case Studies
Figure 8: Examples of segmentation results for error analysis from the (a) FIVES, (b) PolypGen, and (c) ISIC 2018 datasets.
Figure 9: Examples of segmentation results for error analysis from the (a) ATLAS, (b) KiTS23, and (c) TissueNet datasets.
@@
Examples of visualization results of small medical objects in the FIVES, PolypGen, ISIC 2018, ATLAS, KiTS23, and TissueNet datasets are presented in Fig. 8 and Fig. 9. As shown in Fig. 8a, UNet misclassified diabetic retinopathy (green region) as age-related macular degeneration (red region), UNet++, HRNet, & TransNetR misclassified diabetic retinopathy as health retinal vessels (blue region), and PraNet misclassified diabetic retinopathy as glaucoma (yellow region). Besides, HSNet could not detect retinal vasculature in the zoomed-in region. Moreover, none of SOTA methods in the control group can recover retinal vessels on the bottom of the zoomed-in region. Conversely, SvANet not only correctly classified diabetic retinopathy but also effectively detected the retinal vessels’ position and shape.
@@
For the polyps diagnosis as presented in Fig. 8b, SOTA methods such as PraNet either detected a smaller polyp area than the ground truth, or other methods, including HSNet, CFANet, and TransNetR, regarded a larger region than the ground truth. Additionally, UNet, UNet++, and HRNet cannot detect the polyp in the example image. Besides, the detected regions from methods in the control group are far different from the ground truth. However, SvANet not only recognized an area close to the ground truth but also kept similar shapes as ground truth annotations. For the skin lesion examination illustrated in Fig. 8c, methods in the control group caught a larger region that covered the ground truth annotations, bringing the problem that the boundary of skin lesion was underrated. In contrast, SvANet distinguished a skin lesion of a similar size as the ground truth and uncovered its sawtooth-shaped boundary.
@@
For the MRI and CT image modalities analysis, as shown in Fig. 9ab, it is possible to overlook the overlapped medical objects. For instance, UNet++ and PraNet did not discover a tumor in the liver, and HRNet and TransNetR missed a cyst at the edge of the kidney. Besides, UNet and HSNet wrongly recognized a cyst as a tumor. Although the organ region (e.g., liver and kidney) detected by SOTA methods in the control group could be complete, the illness regions, such as the tumor and cyst, were either larger than the ground truth (hepatic tumor) or smaller than the ground truth (cyst) in the example image. However, SvANet correctly sorted out organs and their infected regions. Moreover, SvANet reflected the morphology details of the liver, hepatic tumor, kidney, and cyst in the example image, which are similar to the ground truth annotations.
For tissue cell recognition in the TissueNet dataset, as shown in Fig. 9c, TransNetR and SvANet detailly divided cell boundaries, and they successfully labeled the cells and nuclei regions, as close as resembling the ground truth. In contrast, other SOTA methods struggled to categorize cells and nuclei, and these methods struggled to differentiate cell boundaries, resulting in merging several cells into one.
The visualization results demonstrate SvANet’s capability to differentiate small medical objects across multiple image modalities and object types, consistent with the quantitative results as Sec. IV-B. Thus, SvANet holds tremendous potential for application in the analysis of small medical objects for disease diagnostics, surgeries, etc.
@@
The ablation study presented in Tab. VII demonstrates that using MCAttn exclusively within MCBottleneck delivered at least +0.26% in mDice and +1.06% in sensitivity compared to its application in AssemFormer.
TABLE VII: Impact of the attention modules in MCBottleNeck and AssemFormer. Configurations include (0): None, (1): Squeeze-excitation, and (2) MCAttn. The best results are underlined in bold.
  Attention of Attention of AssemFormer
MCB* inside MCB outside MCB mDice Sensitivity
  (0) (0) (0) 71.81 70.37
(1) (1) (1) 69.99 68.19
(2) (2) (2) 70.54 69.45
(2) (1) (2) 72.57 71.25
(2) (0) (2) 70.92 68.97
(2) (0) (1) 72.32 71.45
(2) (0) (0) 72.58 72.51
(2) (1) (0) 71.49 70.43
(2) (2) (0) 71.82 70.53
         *MCB: MCBottleneck
The Number of Pooled Tensors for MCAttn
The selection of the size and number of pooled tensors for MCAttn is crucial for expanding network variants. We tested combinations (1, 2), (1, 2, 3), (2, 3), and (1, 2, 3, 4). The results, shown in the first, second, and fourth rows of Tab. VIII, reveal that the (1, 2, 3) combination of pooled tensors outperformed (1, 2) and (1, 2, 3, 4) combinations, with improvements exceeding 2.86% & 1.19% in mDice and 4.01% & 1.66%, respectively. Further analysis, as indicated in the second and third rows of Tab. VIII, highlights the necessity of a pool size of 1, leading to an increase of 0.82% in mDice and +2.27% in sensitivity. These findings emphasize the importance of maintaining an optimal level of variation in the network. An insufficient number of pooled tensors can limit performance, whereas an excessive number can introduce too much stochasticity. Thus, striking the right balance is critical for maximizing the effectivness of MCAttn within the model.
TABLE VIII: Size combinations of pooled feature maps in MCAttn (Unit: %). The best results are underlined in bold.
  Pooled tensor sizes mDice Sensitivity
  (1, 2) 69.72 68.50
(1, 2, 3) 72.58 72.51
(2, 3) 71.76 70.24
(1, 2, 3, 4) 71.39 70.85
@@
Given that AssemFormer operates both inside and outside the MCBottleneck, it is essential to explore its impact in these settings. As indicated in Tab. IX, positioning AssemFormer only inside or outside MCBottleneck leads to reductions in performance, with drops ranging from 1.25% $\sim$ 1.30% in mDice and 1.60% $\sim$ 1.86% in sensitivity. Conversely, employing AssemFormer in both locations yields improvements of 0.16% in mDice and 0.67% in sensitivity, enhancing sperms diagnosis. Such results suggest the importance of consistent tensor operations before and after concatenation for cross-scale guidance, as depicted in Fig. 1, ensuring that tensors are processed in the same latent space.
TABLE IX: Impact of the AssemFormer module located inside and outside the MCBottleneck. The best results are underlined in bold.
@@
Inside MCB* Outside MCB mDice Sensitivity
@@
As shown in Fig. 1, post-MCBottleneck concatenation operations can be replaced by using either direct links or addition operations. Nevertheless, the ablation results detailed in Tab. X demonstrate that concatenation for post-MCBottleneck is more effective. Specifically, concatenation delivered improvements of 3.42% & 1.43% in mDice and 6.07% & 3.12% in sensitivity, compared to direct links and addition operations, separately. The findings underscore the importance of preserving feature characteristics and depth information in outputs from the MCBottleneck and cross-scale guidance.
TABLE X: Comparison of concatenation versus addition operations after MCBottleneck. The best results are underlined in bold.

### v4

@@ IV-D1 Main Components Ablation
IV-D3 The Number of Pooled Tensors for MCAttn
The selection of the size and number of pooled tensors for MCAttn is crucial for expanding network variants. We tested combinations (1, 2), (1, 2, 3), (2, 3), and (1, 2, 3, 4). The results, shown in the first, second, and fourth rows of Tab. VI, reveal that the (1, 2, 3) combination of pooled tensors outperformed (1, 2) and (1, 2, 3, 4) combinations, with improvements exceeding 2.86% & 1.19% in mDice and 4.01% & 1.66%, respectively. Further analysis, as indicated in the second and third rows of Tab. VI, highlights the necessity of a pool size of 1, leading to an increase of 0.82% in mDice and +2.27% in sensitivity. These findings emphasize the importance of maintaining an optimal level of variation in the network. An insufficient pooled tensor can limit performance, whereas an excessive number can introduce too much stochasticity. Thus, striking the right balance is critical for maximizing the effectiveness of MCAttn within the model.
TABLE VI: Size combinations of pooled feature maps in MCAttn (Unit: %). The best results are underlined in bold.
  Pooled tensor sizes mDice Sensitivity
  (1, 2) 69.72 68.50
(1, 2, 3) 72.58 72.51
(2, 3) 71.76 70.24
(1, 2, 3, 4) 71.39 70.85
IV-D4 Inference Time
@@
This section quantifies the inference characteristics of the tested networks, including the number of parameters (# Parameters), multiply-accumulate operations (MACs), and inferencing speed. The number of classes was set to eight, and other configurations were consistent with those described in Sec. IV-A2. The inference speed results, averaged over 1000 runs, are presented in Tab. VII. Tab. VII illustrates that SvANet achieved a real-time analysis of medical images with 46.91 frames per second (FPS), which is twice as fast as UNet, 46% faster than nnUNet, and 19% faster than TransNetR. This performance level is suitable for self-examination and clinical diagnostic applications.
TABLE VII: Model size and inference time comparison.
@@
IV-E Negative Case Studies
@@
Examples of visualization results for ultra-small and small medical objects in the FIVES, PolypGen, ISIC 2018, ATLAS, KiTS23, TissueNet, and SpermHealth datasets are presented in Fig. 8 and Fig. 9. As illustrated in Fig. 8a, UNet misclassified diabetic retinopathy (green region) as age-related macular degeneration (red region). Similarly, UNet++, HRNet, nnUNet, and TransNetR misclassified diabetic retinopathy as health retinal vessels (blue region), while PraNet misclassified it as glaucoma (yellow region). Additionally, PraNet struggled to detect retinal vasculature in the zoomed-in region. Furthermore, none of the SOTA methods in the control group could accurately recover the retinal vessels at the bottom of the zoomed-in region. In contrast, SvANet not only correctly classified diabetic retinopathy but also effectively detected the position and shape of retinal vessels.
@@
In the skin lesion examination illustrated in Fig. 8b, UNet, UNet++, HRNet, nnUNet misclassified normal skin as a nevus. Moreover, UNet, PraNet, nnUNet, CFANet, and TransNetR represented the nevus region as a relatively smooth circle, while UNet++ and HRNet captured a larger region encompassing the ground truth annotations, leading to an underestimation of the lesion boundary. In contrast, SvANet accurately identified a skin lesion of similar size to the ground truth and delineated its sawtooth-shaped boundary. For the polyps diagnosis as presented in Fig. 8c, SOTA methods such as PraNet and nnUNet either detected a smaller polyp area than the ground truth, or other methods, including CFANet and TransNetR, regarded a larger region than the ground truth. Besides, UNet, UNet++, and HRNet failed to detect the polyp in the example image. Furthermore, the detected regions from methods in the control group significantly deviated from the ground truth. However, SvANet recognized an area close to the ground truth and maintained shapes akin to ground truth annotations.
Figure 8: Examples of segmentation results across tested methods in the (a) FIVES, (b) ISIC 2018, and (c) PolypGen datasets for error analysis. Examples contain ultra-small objects (i.e., polyp), small objects (i.e., nevus), and objects with $>$10% area ratio (i.e., retinal vessel).
Figure 9: Examples of segmentation results across tested methods in the (a) ATLAS, (b) KiTS23, (c) TissueNet, and (d) SpermHealth datasets for error analysis. Examples contain ultra-small objects (i.e., tumor, cyst, tissue cell, nucleus, and sperm), small objects (i.e., kidney), and objects with $>$10% area ratio (i.e., liver).
@@
For MRI and CT image modalities analysis, as shown in Fig. 9ab, it is possible to overlook the overlapping medical objects, particularly ultra-small ones. For instance, all tested models in the control group failed to identify an ultra-small tumor inside the liver. In addition, HRNet, nnUNet, and TransNetR missed an ultra-small cyst at the edge of the kidney. Moreover, UNet++ and CFANet incorrectly emphasized the background as a tumor or liver region in the example image, and UNet misclassified a cyst as a tumor. Although the organ region (e.g., liver and kidney) detected by SOTA methods in the control group appeared complete, the pathological regions, such as the tumor and cyst, were either larger (hepatic tumor) or smaller (cyst) than the ground truth in the example image. However, SvANet accurately differentiated between organs and their pathological regions. Furthermore, SvANet captured the morphological details of the liver, hepatic tumor, kidney, and cyst in the example image, closely aligning with the ground truth annotations.
For tissue cell recognition in the TissueNet dataset, as shown in Fig. 9c, both TransNetR and SvANet effectively delineated cell boundaries and accurately labeled the cells and nuclei regions, closely resembling the ground truth. In contrast, other SOTA methods struggled to categorize cells and nuclei, leading to difficulties in differentiating cell boundaries and merging several cells into one. For sperm cell analysis, as presented by the final image in Fig. 9d, SvANet precisely located all sperm positions and effectively recognized the region of the broken-line tail of an abnormal sperm. Conversely, tested methods like HRNet struggled to distinguish the broken-line shape. Moreover, UNet, nnUNet, and TransNetR misclassified an unhealthy sperm head as healthy, as indicated by a green subregion in Fig. 9d.
These visualization results align with the findings discussed in Sec. IV-B and Sec. IV-C, suggesting that SvANet holds significant potential for application in general small medical object recognition across various medical imaging modalities for disease diagnostics, surgeries, etc.
V Conclusion
This paper introduces SvANet, a novel network to enhance the segmentation of small medical objects, aiding in the detection of life-threatening diseases and supporting in vitro fertilization. The experimental results demonstrate that the SvANet is significantly effective in distinguishing medical objects of various sizes. SvANet consistently outperformed other SOTA methods, achieving up to 19.95%, 15.03%, 15.01%, 14.64%, 13.57%, 8.09%, and 3.07% increments in mDice for segmenting objects occupying less than 1% image area across TissueNet, FIVES, ISIC 2018, SpermHealth, PolypGen, ATLAS, and KiTS23 datasets. Furthermore, the visualization results confirm that SvANet accurately identifies the locations and morphologies of all medical objects, demonstrating its exceptional capability in segmenting small medical objects. These findings underscore the potential of SvANet as a significant advancement in medical imaging.
Supplemental Materials
Area Computation
The same instance across different images is treated as an independent object because several datasets do not have instance labels during inference. For a fair comparison, each object instance is counted once per image. Area computation results indicate that there are 18% of retinal vasculatures in FIVES [1], 40% of skin lesions in ISIC 2018 [2], 63% of polyps in PolypGen [3], 74% of hepatic tumors and livers in ATLAS [6], 99% of kidneys and corresponding neopasms in KiTS23 [4], and 100% tissue cells in TissueNet [7] under 10% area ratio. The area calculation codes are publically available at https://github.com/anthonyweidai/circle_extractor.
Network Instantiation
@@ V Conclusion
The ablation study presented in Tab. VIII demonstrates that using MCAttn exclusively within MCBottleneck delivered at least +2.04% in mDice and +3.06% in sensitivity compared to its application in AssemFormer.
TABLE VIII: Impact of the attention modules in MCBottleNeck and AssemFormer in the SpermHealth dataset. Configurations include (0): None, (1): Squeeze-excitation, and (2) MCAttn. The best results are underlined in bold.
  Attention in Attention in
MCBottleneck AssemFormer mDice Sensitivity
  (0) (0) 71.81 70.37
(1) (1) 69.99 68.19
(2) (2) 70.54 69.45
(2) (0) 72.58 72.51
@@
Given that AssemFormer operates before and after the concatenation operation, exploring its impact in these settings is essential. As indicated in Tab. IX, positioning AssemFormer only before or after concatenation leads to reductions in performance, with drops ranging from 1.25% $\sim$ 1.30% in mDice and 1.60% $\sim$ 1.86% in sensitivity. Conversely, employing AssemFormer in both locations yields improvements of 0.16% in mDice and 0.67% in sensitivity, enhancing sperms diagnosis. Such results suggest the importance of consistent tensor operations before and after concatenation for cross-scale guidance, as depicted in Fig. 2, ensuring that tensors are processed in the same latent space.
TABLE IX: Impact of the AssemFormer module located before and after concatenation in the SpermHealth dataset. The best results are underlined in bold.
@@
Before After
concatenation concatenation mDice Sensitivity
@@
As shown in Fig. 2, post-MCBottleneck concatenation operations can be replaced using direct links or addition operations. Nevertheless, the ablation results detailed in Tab. X demonstrate that concatenation for post-MCBottleneck is more effective. Specifically, concatenation delivered improvements of 3.42% & 1.43% in mDice and 6.07% & 3.12% in sensitivity, compared to direct links and addition operations, separately. The findings underscore the importance of preserving feature characteristics and depth information in outputs from the MCBottleneck and cross-scale guidance.
TABLE X: Comparison of concatenation versus addition operations after MCBottleneck in the SpermHealth dataset. The best results are underlined in bold.